# This work is supported by the ML4Seismic Industry Partners at Georgia Tech

### v1

@@ V-A3 Model Capacity and Transferability
A factor that also plays a role in modulating generalization is model capacity. Fig. 12 and Fig. 13 show that large models like segformer benefit from pretraining on CRACKS, while smaller models (e.g., hed, deeplab-m) generalize better when trained from scratch. The observation suggests that both data alignment and model capacity affect transfer effectiveness.
Furthermore, models respond differently to fine-tuning. When pretraining on FaultSeg3D and fine-tuning on Thebe, models degrade in CRACKS performance. Fig. 14) shows that Thebe induces domain shifts that are hard to unlearn. Conversely, even though the finetuning dataset is distributionally closer to the target, Thebe-pretrained segformer and deeplab also degrade on CRACKS after finetuning on FaultSeg3D as shown in Fig.  14. The asymmetric behavior highlights the difficulty of finding universally robust pretraining strategies.
@@
Figure 13: Individual models tested on Thebe. (a) Models pre-trained on the FaultSeg3D data, and fine-tuned on different data. (b) Models pre-trained on the CRACKS data, and fine-tuned on different data.
@@ V-B1 Window Size and Loss Function
We evaluate the unet model with a ResNet50 backbone using Dice and Binary Cross-Entropy (BCE) losses, as well as across different window sizes: 96, 128, 256 and 512, or up to the size allowed by the original sections in each dataset. The results for these experiments are shown in Fig. 15, where bigger markers correspond to bigger patch sizes. We can see that when using Dice as a loss, models benefit from having access to larger window sizes, while the same does not hold when using the BCE loss. This discrepancy is due to the class imbalance nature of the fault delineation task: BCE tends to work best when foreground and background are balanced but fails otherwise, and Dice is generally designed for class-imbalanced scenarios [116]. In this context, using smaller patches is equivalent to ’zooming into’ the small fault regions and therefore reducing the class imbalance, thus improving BCE performance. Overall, this further validates our choice of using the Dice loss on our benchmarking experiments.
@@ V-B2 Model Nuances in Fault Delineation
We also qualitatively observe that each of the evaluated models presents different nuances in the structure of their fault predictions, irrespective of the pretraining or finetuning strategy used, many of which can be observed in Figs. 19 and 20. For example, deeplab tends to produce irregular, stair-like faults, while segformer produces thicker, blob-like faults. unet architectures in general tend to produce thinner faults, with unet++ generating more fragmented ones. These architectural signatures are consistent across training setups and highlight the influence model design choices inherently have in shaping the morphology of predicted faults, which is an important consideration when selecting models for downstream tasks or when interpreting evaluation results beyond numerical metrics.
@@ V-C1 Sensitivity to Visual Structure
Distance-based metrics are generally more tolerant to the structure of the predicted fault. This behavior is illustrated in Fig.  LABEL:fig:dicevsbcd. Where Fig. LABEL:fig:dicevsbcd_a represents the ground truth, while LABEL:fig:dicevsbcd_b and LABEL:fig:dicevsbcd_c show the predictions of two different models. Although the prediction in Fig. LABEL:fig:dicevsbcd_b appears structurally closer to the ground truth, it receives significantly worse BCD and Hausdorff scores compared to the prediction in Fig. LABEL:fig:dicevsbcd_c. Since distance-based metrics do not heavily weigh continuity, the inclusion of a few extra pixels around the fault can improve the BCD score even if those pixels lack proper structural alignment. Notably, these patterns are not anomalies, different models often generate such consistent outputs. Consequently, numerical metrics can be misleading and may display discrepancies between one another.
@@ V-C2 Fault Sparsity
Another issue with distance-based metrics is that they are designed to measure the quality of a single continuous object in the image, whereas faults can consist of multiple sparse objects. This makes distance-based metrics very sensitive to both the number of faults present in the ground truth and their sparsity. We showcase this issue in Fig.  17, where we consider two cases: one with a few faults and another with many faults. Noise is added to both images, and each is compared against its original version. The case with fewer faults contains many outliers compared to the case with many faults, since the added noise often lies far from any existing fault. As demonstrated in Fig.  4, because BCD is bidirectional, it accounts for each added noise pixel by searching for its closest existing fault. Therefore, having fewer faults results in a much worse BCD score. DICE, while also penalizing noise, is more stable across different sparsity levels.
@@ V-C3 Contradictory Scores and Human Judgement
Although pixel-based metrics are more stable, they remain sensitive to slight pixel shifts. In Fig.  18, we show examples corresponding to two different models. The prediction in Fig. 18(c) looks visually closer to the ground truth, but receives a worse Dice score (0.1325 vs. 0.13421) and significantly better BCD and Hausdorff (40.782 vs. 115.436 and 29.0 vs 72.5). Since the faults in both predictions are poorly structured and spatially misaligned, the Dice coefficient (being overlap-based) penalizes them similarly. In contrast, distance-based metrics such as BCD and Hausdorff distances are more sensitive to the spatial coherence of the predicted faults, hence, they are more tolerant to the structure of the faults. These contradictions imply that some metrics may conflict with visual intuition or downstream utility, and point to the need for context-aware metric selection frameworks.
@@
Figure 20: Predictions of the models pretrained on CRACKS tested Thebe
TABLE II: Key observations
@@
VI Conclusion
In this work, we present the first large-scale benchmarking study of machine learning models for seismic fault delineation under domain shift. By systematically training and evaluating over 200 models across three diverse seismic datasets, we provide a comprehensive analysis of generalization, transferability, and evaluation challenges in seismic segmentation. Our results reveal the fragility of fine-tuning strategies when faced with shifts in data distribution, the emergence of catastrophic forgetting when moving across domains, and the sensitivity of model behavior to architecture and pretraining alignment. We also highlight the limitations of standard evaluation metrics, which often fail to capture geological plausibility, and we advocate for the development of more nuanced frameworks that account for expert ambiguity and structural realism. Collectively, these findings establish a robust empirical foundation for understanding the tradeoffs in current seismic ML workflows and offer practical guidance to researchers and practitioners seeking to build models that are not only accurate but also transferable, interpretable, and aligned with real-world exploration needs.

### v2

@@ V-A3 Model Capacity and Transferability
A factor that also plays a role in modulating generalization is model capacity. Fig. 8 shows that large models like segformer benefit from pretraining on CRACKS, while smaller models (e.g., hed, deeplab-m) generalize better when trained from scratch. The observation suggests that both data alignment and model capacity affect transfer effectiveness. This interaction between model size and pretraining benefits suggests that a higher model capacity or parameter count is not always better, but rather amplifies both the advantages and disadvantages of transfer. Large-capacity architectures can better exploit well-aligned pretrained features but are also more prone to overfitting to domain-specific artifacts when source and target distributions diverge. Smaller-capacity models, by contrast, may act as an implicit regularizer, limiting over-specialization and enabling better zero-shot transfer in mismatched scenarios, as can be observed in Fig. 8.
Furthermore, models respond differently to fine-tuning. When pretraining on FaultSeg3D and fine-tuning on Thebe, models degrade in CRACKS performance (Fig. 9, blue circle) shows that Thebe induces domain shifts that are hard to unlearn. Conversely, even though the finetuning dataset is distributionally closer to the target, Thebe-pretrained segformer and deeplab also degrade on CRACKS after finetuning on FaultSeg3D as shown in Fig.  9 (green circles). The asymmetric behavior highlights the difficulty of finding universally robust pretraining strategies.
[…]
TABLE II: Results on domain adaptation case study
Training configuration FDA DANN
Dice Hausdorff Dice Hausdorff
FaultSeg3D $\rightarrow$ Cracks 0.139 48.221 0.156 41.058
FaultSeg3D $\rightarrow$ Thebe 0.276 20.277 0.319 15.872
@@ V-B1 Window Size and Loss Function
We evaluate the unet model with a ResNet50 backbone using Dice and Binary Cross-Entropy (BCE) losses, as well as across different window sizes: 96, 128, 256 and 512, or up to the size allowed by the original sections in each dataset. The results for these experiments are shown in Fig. 10, where bigger markers correspond to bigger patch sizes.
We can see that two trends emerge in these experiments. First, when using Dice as a loss, models benefit consistently from larger window sizes, with performance generally improving as the spatial context grows. This is likely because larger patches provide the network with a more complete view of fault structures, enabling better continuity modeling across sections. Second, when using BCE loss, this trend does not hold: performance remains flat or slightly declines with larger windows. This discrepancy stems from the class imbalance inherent to fault delineation [140]. BCE tends to work best when foreground and background classes are more balanced, whereas Dice loss is explicitly designed to handle imbalance. Using smaller patches effectively “zooms in” on the sparse fault regions, increasing the proportion of fault pixels and improving BCE performance.
From a practical perspective, these results suggest that in real-world seismic interpretation, the optimal windowing standard depends on the chosen loss function and the model’s need for contextual information. For losses that handle imbalance well (e.g., Dice), larger tiles are preferable because they capture longer fault segments and contextual cues that allow for improve spatial continuity. For imbalance-sensitive losses (e.g., BCE), smaller tiles may sometimes be beneficial, though at the cost of reduced global context. Given that Dice consistently benefits from larger windows, and that most modern segmentation pipelines for faults employ class-imbalance aware losses, our choice of using Dice with the largest possible window sizes in our benchmark experiments is both empirically supported and aligned with best practices in seismic fault delineation.
@@ V-B2 Model Nuances in Fault Delineation
We also qualitatively observe that each of the evaluated models presents different nuances in the structure of their fault predictions, irrespective of the pretraining or finetuning strategy used, many of which can be observed in Fig. 14. For example, deeplab tends to produce irregular, stair-like faults, while segformer produces thicker, blob-like faults. unet architectures in general tend to produce thinner faults, with unet++ generating more fragmented ones. These architectural signatures are consistent across training setups and highlight the influence model design choices inherently have in shaping the morphology of predicted faults, which is an important consideration when selecting models for downstream tasks or when interpreting evaluation results beyond numerical metrics.
@@ V-C1 Sensitivity to Visual Structure
Distance-based metrics are generally more tolerant to the structure of the predicted fault. This behavior is illustrated in Fig.  11. Where Fig. 11(a) represents the ground truth, while 11(b) and 11(c) show the predictions of two different models. Although the prediction in Fig. 11(b) appears structurally closer to the ground truth, it receives significantly worse BCD and Hausdorff scores compared to the prediction in Fig. 11(c). Since distance-based metrics do not heavily weigh continuity, the inclusion of a few extra pixels around the fault can improve the BCD score even if those pixels lack proper structural alignment. Notably, these patterns are not anomalies, different models often generate such consistent outputs. Consequently, numerical metrics can be misleading and may display discrepancies between one another.
@@ V-C2 Fault Sparsity
Another issue with distance-based metrics is that they are designed to measure the quality of a single continuous object in the image, whereas faults can consist of multiple sparse objects. This makes distance-based metrics very sensitive to both the number of faults present in the ground truth and their sparsity. We showcase this issue in Fig.  12, where we consider two cases: one with a few faults and another with many faults. Noise is added to both images, and each is compared against its original version. The case with fewer faults contains many outliers compared to the case with many faults, since the added noise often lies far from any existing fault. As discussed in Section II-E, because BCD is bidirectional, it accounts for each added noise pixel by searching for its closest existing fault. Therefore, having fewer faults results in a much worse BCD score. DICE, while also penalizing noise, is more stable across different sparsity levels.
@@ V-C3 Contradictory Scores and Human Judgement
Although pixel-based metrics are more stable, they remain sensitive to slight pixel shifts. In Fig.  13, we show examples corresponding to two different models. The prediction in Fig. 13(c) looks visually closer to the ground truth, but receives a worse Dice score (0.1325 vs. 0.13421) and significantly better BCD and Hausdorff (40.782 vs. 115.436 and 29.0 vs 72.5). Since the faults in both predictions are poorly structured and spatially misaligned, the Dice coefficient (being overlap-based) penalizes them similarly. In contrast, distance-based metrics such as BCD and Hausdorff distances are more sensitive to the spatial coherence of the predicted faults, hence, they are more tolerant to the structure of the faults. These contradictions imply that some metrics may conflict with visual intuition or downstream utility, and point to the need for context-aware metric selection frameworks.
@@
VI Analysis Using Fault Characteristic Metrics
TABLE III: Summary statistics of the three datasets: CRACKS (A), Thebe (B), and FS (C). Values are reported as mean $\pm$ standard deviation.
Dataset Length Curvature Sinuosity Segments Stepover Density
CRACKS 6577.75 $\pm$ 1090.82 0.0177 $\pm$ 0.0038 95.35 $\pm$ 54.30 47.22 $\pm$ 7.61 0.978 $\pm$ 0.0046
Thebe 38457.25 $\pm$ 7676.67 0.0261 $\pm$ 0.0058 13.46 $\pm$ 12.32 10.45 $\pm$ 1.85 0.902 $\pm$ 0.0169
FS 16543.79 $\pm$ 3452.73 0.0369 $\pm$ 0.0081 5.70 $\pm$ 6.32 3.18 $\pm$ 1.10 0.623 $\pm$ 0.207
TABLE IV: Evaluation metrics when models are pretrained on Thebe and tested on Thebe, with or without finetuning on other datasets.
Setup Strike Sim. Curvature Curv. RMSE Curv. Corr Sinuosity $\Delta$ Sinuosity Ratio Length $\Delta$ Length Ratio Segment $\Delta$ Segment Ratio Stepover $\Delta$ Stepover Ratio
Thebe Only 0.907 $\pm$ 0.081 -0.001 $\pm$ 0.006 0.128 $\pm$ 0.014 0.019 $\pm$ 0.054 0.246 $\pm$ 0.140 0.829 $\pm$ 0.068 1715.41 $\pm$ 1044 1.649 $\pm$ 0.543 51.47 $\pm$ 10.04 3.292 $\pm$ 0.455 1.779 $\pm$ 6.669 0.993 $\pm$ 0.061
Thebe→CRACKS 0.756 $\pm$ 0.107 0.012 $\pm$ 0.010 0.174 $\pm$ 0.015 -0.010 $\pm$ 0.044 0.080 $\pm$ 0.152 0.950 $\pm$ 0.090 -1593.73 $\pm$ 925.87 0.560 $\pm$ 0.187 67.08 $\pm$ 9.09 4.001 $\pm$ 0.490 4.770 $\pm$ 0.001 0.989 $\pm$ 0.104
Thebe→FS 0.658 $\pm$ 0.053 0.007 $\pm$ 0.010 0.166 $\pm$ 0.017 -0.003 $\pm$ 0.049 -0.055 $\pm$ 0.183 0.968 $\pm$ 0.114 1141.05 $\pm$ 963.78 0.705 $\pm$ 0.241 90.39 $\pm$ 14.91 5.046 $\pm$ 0.788 0.0004 $\pm$ 0.0001 1.004 $\pm$ 0.024
When evaluating transfer setups where all models were pretrained on Thebe and subsequently finetuned on different datasets before being tested back on Thebe, we observe a clear correspondence between the statistical characteristics of the finetuning dataset and the error patterns in the predictions. Training and testing solely on Thebe produces the strongest alignment with the ground truth, as indicated by the highest strike similarity (0.907), reflecting that the model effectively captures the long and continuous fault structures characteristic of Thebe. However, this setup also leads to an overestimation of fault length (length ratio of 1.65) and moderate over-segmentation (segment ratio of 3.29), suggesting that the model tends to exaggerate continuity while artificially fragmenting longer structures.
In contrast, when the Thebe-pretrained model is finetuned on CRACKS and evaluated on Thebe, performance degrades in ways consistent with the CRACKS dataset statistics, which emphasize short, jagged, highly segmented faults with frequent crossovers. Strike similarity drops substantially to 0.756, indicating that the orientation of predictions becomes less aligned with Thebe’s long faults. Length is strongly underestimated (length ratio of 0.56), directly mirroring the shorter average fault lengths in CRACKS. Segmentation increases (segment ratio of 4.00), and the stepover delta rises to 4.77, reflecting the fragmentation and crossover-heavy nature of CRACKS. Moreover, spacing becomes strongly negative, showing that the predictions adopt the denser fault placement bias of CRACKS rather than the sparser spacing of Thebe.
A different type of degradation is observed when finetuning on FaultSeg3D, which is dominated by medium-length, smooth faults with very low segmentation and reduced connectivity. Here, strike similarity falls further to 0.658, the lowest among all setups, demonstrating poor alignment with Thebe’s long continuous faults. Length is again underestimated (ratio of 0.70), in line with FaultSeg3D’s shorter structures. Segmentation is highly exaggerated (segment ratio of 5.05), but stepovers almost vanish (stepover delta close to 0), consistent with FaultSeg3D’s low stepover density. The negative spacing indicates that the model no longer respects Thebe’s distribution of fault separations. Thus, while finetuning on FaultSeg3D removes the crossover effects seen with CRACKS, it introduces extreme fragmentation of Thebe’s continuous faults and produces disconnected structures.
In summary, the in-domain Thebe-only model best preserves structural alignment but exaggerates continuity and segmentation, whereas finetuning on CRACKS injects a bias toward jagged, fragmented, and crossover-rich structures, and finetuning on FaultSeg3D enforces smooth, disconnected, and overly fragmented representations. Both transfer setups degrade performance relative to the in-domain baseline, but in ways that directly reflect the statistical properties of the finetuning datasets. This demonstrates that the dataset-specific fault geometry strongly governs the inductive bias of models, even when pretrained on the same source.
TABLE V: Key observations
@@ VI Analysis Using Fault Characteristic Metrics
Domain adaptation FDA and DANN improve large-shift transfers (i.e. FaultSeg3D to Thebe) but degrade performance in aligned domains (i.e. FaultSeg3D to CRACKS).
@@
Loss–context relationship Dice loss benefits from larger window sizes (captures fault continuity), while BCE is more effective with smaller patches due to class imbalance.
Fault characteristic transfer Models inherit structural/geometric biases of finetuning dataset. For instance, CRACKS fragmented and crossover faults, while FaultSeg3D induces smoother and disconnected faults.
VII Conclusion
In this work, we present the first large-scale benchmarking study of seismic fault delineation models under domain shift, spanning more than 200 training configurations across three heterogeneous datasets. Our results show that fine-tuning is generally effective when source and target domains are closely aligned, but becomes brittle under stronger shifts, often leading to catastrophic forgetting. Model capacity also modulates transferability: larger architectures such as Segformer tend to adapt more effectively, while smaller models are more sensitive to mismatch. Domain adaptation methods such as FDA and DANN proved beneficial in highly divergent transfers but sometimes degraded performance in more similar settings, highlighting the risk of negative transfer.
Beyond conventional metrics such as Dice or Hausdorff distance, our geometric and topological analysis demonstrated that models also absorb the structural biases of the datasets they are finetuned on: CRACKS-trained models tended to reproduce short, jagged, crossover-rich structures, while FaultSeg3D-trained models favored smoother but disconnected faults. Even when pretrained on the same source, prediction styles were shaped by the statistical properties of the finetuning dataset. These findings underscore that evaluation should not rely solely on pixel-level accuracy but must also account for structural plausibility and geological realism.
At the same time, important open questions remain. Future research should focus on designing adaptation methods that remain effective across both mild and severe domain shifts, incorporating geological priors into model architectures to mitigate dataset-specific biases, and developing evaluation protocols that balance quantitative rigor with geological interpretability. Addressing these challenges will be essential to building seismic DL pipelines that are not only accurate but also dependable in real-world interpretation workflows.